The great circle has an area of 289pie square feet.
Calculate the surface area and volume.

Answer by jim_thompson5910(35256)

A = pi*r^2

289pi = pi*r^2

289 = r^2

r^2 = 289

r = sqrt(289)

r = 17

The radius of this sphere is 17 feet.

-------------------------------------------------------

Surface Area:

SA = 4pi*r^2

SA = 4pi*17^2

SA = 1156pi

The exact surface area of this sphere is 1156pi

-------------------------------------------------------

Volume:

V = (4/3)pi*r^3

V = (4/3)pi*17^3

V = (19652/3)pi

The exact volume of this sphere is (19652/3)pi